OpenStudy (anonymous):

WHAT IS 1/2h+1/3(h-6)=5/6h+2

OpenStudy (anonymous):

\(\dfrac{1}{2}h+\dfrac{1}{3}(h-6)=\dfrac{5}{6}h + 2\) Order of operations: Parenthesis - distribute first.

OpenStudy (anonymous):

I would use cross multiplication.

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

3(h-6) can be put on top of the fraction due to fraction multiplication rules.

OpenStudy (anonymous):

oh I remember okay I got it!

OpenStudy (anonymous):

so it becomes 3(h-6)/2h+1

OpenStudy (anonymous):

= the other mess

OpenStudy (anonymous):

-4 = 0 is the answer there is no solution :)

OpenStudy (anonymous):

you can check your answer using wolframalpha